1. What is your assigned network address? 224.151.73.0/24

2. What is the subnet mask (dotted decimal form) for this network?

3. What is the broadcast address for this network?

4. How many addressable hosts are available for this network?

Following good network design practices, you decide to segment this network into 4 equal subnets, to be used or shared among the different departments: Software Development, Technical Support, Marketing/Sales, and Management & Human Resources.

5. To create 4 subnets, you must borrow how many bits from the host portion of the network? (Hint: solve 4 = 2n)

6. What is the subnet mask (dotted decimal form) for these new subnets?
Answer #1

1. What is your assigned network address? 224.151.73.0/24

Solution:

Network Address is: 224.151.73.0

calculated as AND of given IP (224.151.73.0) with its MASK (255.255.255.0) give network address

2. What is the subnet mask (dotted decimal form) for this network?

Solution:

255.255.255.0 (/24 shows first 24 bits are 1)

3. What is the broadcast address for this network?

Solution:

224.151.73.255 (This is the last address of subnet)

4. How many addressable hosts are available for this network?

Solution:

28 - 2 =256 - 2 = 254

The address range is 224.151.73.1 - 224.151.73.254

Following good network design practices, you decide to segment this network into 4 equal subnets, to be used or shared among the different departments: Software Development, Technical Support, Marketing/Sales, and Management & Human Resources.

5. To create 4 subnets, you must borrow how many bits from the host portion of the network? (Hint: solve 4 = 2n)

Solution:

as 4 subnets and hence 4 bits borrow from the host.

6. What is the subnet mask (dotted decimal form) for these new subnets?

Solution:

255.255.255.192

Subnet Name Address Dec Mask Assignable Range Broadcast
Subnet 1 224.151.73.0 255.255.255.192 224.151.73.1 - 224.151.73.62 224.151.73.63
Subnet 2 224.151.73.64 255.255.255.192 224.151.73.65 - 224.151.73.126 224.151.73.127
Subnet 3 224.151.73.128 255.255.255.192 224.151.73.129 - 224.151.73.190 224.151.73.191
Subnet 4 224.151.73.192 255.255.255.192 224.151.73.193 - 224.151.73.254
224.151.73.255